Show that

(

)

(

)

(

)

1,

,

1

,

CnkCnkCnk

+ =

− +

.



1
Expert's answer
2021-10-18T13:36:12-0400

Show that "C(n,k)=C(n,n-k)" :


"c(n,k)=\frac{n!}{k!(n-k)!}=\frac{n!}{(n-k)!(n-(n-k))!}=C(n,n-k)"